When each side of a given square is increased by 4 feet the are is increased by 64 square feet. Determine the dimensions of the original square.
Answer by shweta(56)

Let the side of a square be 's'feet
Area of the square= s*s=s^2 feet
Now side increases by 4 feet, s+ 4
The area also increases by 64 feet
We get,
(s+4)^2= s^2 + 64 ...(1)
Apply the formula (a+b)^2= a^2 +b^2+ 2*a*b
s^2 + 4^2+ 2*s*4= s^2+ 64
s^2 on both the sides get cancelled
4^2+ 8s= 64
16+ 8s= 64
We have to shift the variable on one side and the numbers on the other side of 'equal to'sign
8s= 64- 16
8s= 48
s= 48/8 ( here 48 is divided by 8 because on the other side of 'equal to' sign 8 was multiplied to s)
s=6